# Geometry-informed Channel Statistics Prediction Based upon Uncalibrated Digital Twins thanks: This research is supported by the HORIZON JU-SNS-2022-STREAM-B-01-02 CENTRIC project (Grant Agreement No.101096379) and the Villum Investigator Grant “WATER” from the Velux Foundations, Denmark. Authors Mahmoud Saad Abouamer and Robin J. Williams contributed equally to this work

### v1

@@ Abstract
Digital twins (DTs) of wireless environments can be utilized to predict the propagation channel and reduce the overhead of required to estimate the channel statistics. However, direct channel prediction requires data-intensive calibration of the DT to capture the environment properties relevant for propagation of electromagnetic signals. We introduce a framework that starts from a satellite image of the environment to produce an uncalibrated DT, which has no or imprecise information about the materials and their electromagnetic properties. The key idea is to use the uncalibrated DT to implicitly provide a geometric prior for the environment. This is utilized to inform a Gaussian process (GP), which permits the use of few channel measurements to attain an accurate prediction of the channel statistics. Additionally, the framework is able to quantify the uncertainty in channel statistics prediction and select rate in ultra-reliable low-latency communication (URLLC) that complies with statistical guarantees. The efficacy of the proposed geometry-informed GP is validated using experimental data obtained through a measurement campaign. Furthermore, the proposed prediction framework is shown to provide significant improvements compared to the benchmarks where i) direct channel statistics prediction is obtained using an uncalibrated DT and (ii) the GP predicts channel statistics using information about the location.
@@ I Introduction
Digital twins (DTs) that accurately model wireless channels have demonstrated significant utility for both communication and localization tasks [1]. For instance, by maintaining a precise DT, site-specific digital twins can be used to reduce CSI acquisition and feedback overhead [2]. Additionally, DT-generated fingerprinting databases be used to improve positioning accuracy [3]. However, for the DT to provide accurate predictions, its electromagnetic properties should match real-world channel conditions. The process of configuring the objects and materials in the DT to represent the wireless environment is referred to as calibration.
@@
Differentiable ray tracers (RT), such as Sionna RT [4] are instrumental in enabling effective calibration methods. Differentiable RTs allow the computation of gradients for system parameters, enabling tasks like data-driven material property optimization. This was demonstrated in recent works, such as [5, 6], which leverage differentiable RT for material optimization. For instance, to address phase errors caused by geometric discrepancies, [6] employs a variational expectation maximization algorithm for mitigating phase errors and thereby improving prediction accuracy.
This work treats the problem of resource allocation for ultra reliable low latency communications. The problem is to pick a data-rate which satisfies a set of reliability and latency constraints. Due to the latency constraint, channel feedback procedures are not always feasible. In such cases, the system must rely on alternatives such as channel statistics maps to select a rate which satisfies the reliability constraint with a sufficiently high probability. Indeed, a fully calibrated DT which accurately represents the objects and materials in a scene, enables the direct prediction of channel statistics. However, the process of frequently calibrating a DT to achieve this level of accuracy is resource-intensive, restricting its practical use to limited scenarios.
@@
In contrast with prior works, this work presents a novel method that utilizes uncalibrated DTs generated from satellite images obtained from open-source maps, such as OpenStreetMap [7]. Rather than relying on a perfectly calibrated DT, the uncalibrated DT is employed to extract geometric features of the environment. These geometric features are combined with sparse channel measurements taken from a few anchor points to estimate the channel statistics across the entire scene. To enables this, we leverage the spatial interpolation capabilities of a Gaussian process (GP) along with the geometric consistency of the DT to predict channel statistics with high accuracy. Thus, by providing a geometry-informed GP framework for channel prediction, this work aims to improve the overall reliability and performance of wireless systems in a more efficient and scalable manner. The results show that the proposed geometry-informed GP significantly improves prediction accuracy across the entire scene compared to the benchmarks where (i) an uncalibrated DT directly predicts channel statistics and (ii) a GP uses location information to predict channel statistics. Moreover, when the channel statistics map is used in rate selection for URLLC, the proposed framework is shown to (approximately) double the data rate while maintaining the same reliability, compared to the benchmarks.
The paper is organized as follows. Sec. II describes the system model. Sec. III describes the scenario upon which the proposed framework is evaluated as well as the construction of the DT. Sec. IV describes the proposed methods for estimating the channel statistics map. Sec. V presents a range of simulation results based on the measured data. Finally, sec. VI concludes this paper.
Notation: $\mathrm{i}\mkern 1.0mu$ is the imaginary unit, $||\absolutevalue{\cdot}$ is the absolute value, vectors are denoted by bold lowercase symbols, $\mathbb{E}[x]$ is the expectation operator.
II System model
This work considers a single receiver and a single transmitter, both equipped with a single antenna each. The transmitter is a assumed to have access to an estimate of its location. The location of the receiver is static and known to all parties. Assuming an Orthogonal Frequency Division Multiplexing (OFDM) system, perfect clock synchronization, a channel without Doppler shifts, and delay-spread equalization, such as cyclic-prefixing, the signal received at the receiver at a given subcarrier is:
$y=\sqrt{P_{\text{tx}}}h\left(\mathbf{x}\right)s+n$ (1)
@@ II System model
where $h\left(\mathbf{x}\right)\in\mathbb{C}$ is the frequency domain channel between the transmitter and receiver, $\mathbf{x}\in\mathbb{R}^{3\times 1}$ is the transmitter location, $s\in\mathbb{C}$ is a random transmit symbol with mean $\mu_{s}=\mathbb{E}[s]=0$ and variance $\sigma_{s}^{2}=\mathbb{E}[ss^{\mathsf{H}}]=1$, $P_{\text{tx}}\in\mathbb{R}$ is the time-averaged transmit power, and $n\sim\mathcal{CN}\left(0,\sigma_{n}^{2}\right)$ is additive white Gaussian noise. The time-averaged signal to noise ratio is thus
[…]
The maximum rate that the channel can sustain under error-free communications is given through the channel capacity as [8]
$R_{\text{max}}=B\cdot\log_{2}\left(1+\gamma\left(\mathbf{x}\right)\right)$ (3)
where $B$ is the bandwidth of the subcarrier. Given known channel state information, the rate can thus be chosen in a straight forward manor as $R=R_{\text{max}}$. The channel coefficient, $h\left(\mathbf{x}\right)$, is dependent on the propagation environment and it is commonly modeled as a realization of a random distribution. The randomness is in part due to imperfect channel estimation, dynamic changes of the environment, and inaccurate position estimation. The DT can be used to remove some of the uncertainty from $h\left(\mathbf{x}\right)$, which, in effect, changes the distribution of $h\left(\mathbf{x}\right)$ from the perspective of a system designer. The problem is then to find the maximum rate $R$ for which the probability of outage, $p_{\text{out}}$, is less than $\epsilon$;
$\epsilon>p_{\text{out}}\left(R\right)$ $=P\left(R_{\text{max}}<R\right),$
$=P\left(B\cdot\log_{2}\left(1+\frac{P_{\text{tx}}\absolutevalue{h\left(\mathbf{x}\right)}^{2}}{\sigma_{n}^{2}}\right)<R\right),$
@@
$\implies R<B\cdot\log_{2}\left(1+\frac{P_{\text{tx}}p_{\epsilon}\left(\mathbf{x}\right)}{\sigma_{n}^{2}}\right)$ (4)
where $P\left(a<b\right)$ is the probability that $a<b$ and $p_{\epsilon}\left(\mathbf{x}\right)$ is the $\epsilon$-quantile the channel power which ideally can be obtained from the of the cumulative probability distribution of $p\left(\mathbf{x}\right)=\absolutevalue{h\left(\mathbf{x}\right)}^{2}$, given that the distribution of $h\left(\mathbf{x}\right)$ is known. However, due to latency constants, just as the instantaneous value of $h\left(\mathbf{x}\right)$ is unknown, so is the distribution of $h\left(\mathbf{x}\right)$.
Assuming, similarly to [9], that the the logarithm of the $\epsilon$-quantile channel power $q_{\epsilon}\left(\mathbf{x}\right)=\ln\left(p_{\epsilon}\left(\mathbf{x}\right)\right)$ behaves as a multivariate Gaussian random variable across space, the power quantiles can be predicted over space through Gaussian process regression, given a set of known power quantiles. The estimation is done on the logarithm of the power quantiles as the logarithm provides a more Gaussian-like process. Consider the widely applied wide sense stationary uncorrelated scattering (WSSUS) channel model, in which the channel power, $p\left(\mathbf{x}\right)\sim\lambda\left(\mathbf{x}\right)\mathrm{e}^{-\lambda\left(\mathbf{x}\right)p}\,|\,p\geq 0$, follows an exponential distribution [10]. The natural logarithm of the channel power, $q\left(\mathbf{x}\right)=\ln\left(p\left(\mathbf{x}\right)\right)\sim\lambda\left(\mathbf{x}\right)\mathrm{e}^{q-\lambda\left(\mathbf{x}\right)\mathrm{e}^{q}}$, has a more Gaussian-like distribution and is thus better approximated by a multivariate Gaussian.
Obtaining spatial samples of the power quantiles, $p_{\epsilon}\left(\mathbf{x}\right)$, necessary for ultra reliable communications, upon which to predict through Gaussian process regression, requires either relying on model based approaches which are susceptible to model mismatch errors, or a very high number of channel measurements to establish an empirical distribution, which results in high latency. An alternative approach proposed in [11] is to a range of subcarriers and use frequency as a proxy for space. By measuring a wide bandwidth, an empirical channel power distribution can be established and an estimate of the $\epsilon$-quantile of the channel power, $p_{\epsilon}\left(\mathbf{x}\right)$, can be extracted.
The validity of using frequency as a proxy for space can be argued by considering again the WSSUS Rayleigh fading channel [10]. In the farfield of the transmitter, the impinging field can be written as a position dependent distribution of planewaves. In quantizing the distribution, the received signal in the vicinity of a point $\boldsymbol{x}$ can be written as
$h\left(\mathbf{x},f\right)\approx s\sum_{n=1}^{N}\alpha_{n}\mathrm{e}^{-\mathrm{i}\mkern 1.0muk\mathbf{x}^{\mathsf{T}}\mathbf{r}_{n}}\mathrm{e}^{-\mathrm{i}\mkern 1.0mu\tau_{n}f}$ (5)
where $k\in\mathbb{R}$ is the wavenumber, $\mathbf{x}\in\mathbb{R}^{3\times 1}$ is a point in the vicinity of $\boldsymbol{x}$ and $\alpha_{n}\in\mathbb{C}$ is the amplitude and phase, $\mathbf{r}_{n}\in\left\{\mathbb{R}^{3\times 1}|\mathbf{r}_{n}^{\mathsf{T}}\mathbf{r}_{n}=1\right\}$ is the normal, and $\tau_{n}$ is the excess delay of the $n$’th planewave. In the Rayleigh channel, $\alpha_{n}$ is a set of zero mean independent identically distributed circular symmetric complex random variables. Independence between all $\tau_{n}$, $\mathbf{r}_{n}$, and $\alpha_{n}$ is assumed. Due to $\alpha_{n}$ being circular symmetric complex random variables, $\mathbb{E}[h\left(\mathbf{x},f\right)]=0$. Additionally, due to the central limit theorem, $h\left(\mathbf{x},f\right)\sim\mathcal{CN}\left(0,\sigma^{2}\right)$ for high $N$, where $\sigma^{2}=\mathbb{E}\left[\absolutevalue{h\left(\mathbf{x},f\right)}^{2}\right]$. To allow relating the power quantile obtained from sampling across frequency to the power quantile obtained from sampling across space,
@@
Figure 1: Two pictures of the scenario. Left: A satellite picture of the outdoor scenario not taken at the day of the measurement. Transmitter locations marked with blue dots [11]. Right: Illustration of the geometry made using the sionna.rt.scene.render python library [12] and data extracted from OpenStreetMap [7]. Transmitter points are marked as green dots.
the covariance properties of the channel across frequency and space, is analyzed. The channel covariance function is given as
$\Sigma^{2}\left(\mathbf{x}_{1},\mathbf{x}_{2},f_{1},f_{2}\right)$ $=\mathbb{E}\left[h\left(\mathbf{x}_{1},f_{1}\right)h^{*}\left(\mathbf{x}_{2},f_{2}\right)\right],$
=|$=\absolutevalue{s}^{2}\sigma^{2}c_{\mathbf{x}}\left(\mathbf{x}_{1}-\mathbf{x}_{2}\right)c_{f}\left(f_{1}-f_{2}\right),$ (6)
where $c_{\mathbf{x}}\left(\mathbf{x}\right)=\mathbb{E}\left[\mathrm{e}^{-\mathrm{i}\mkern 1.0muk\mathbf{x}^{\mathsf{T}}\mathbf{r}_{n}}\right]$ is the spatial correlation function, and $c_{f}\left(f\right)=\mathbb{E}\left[\mathrm{e}^{-\mathrm{i}\mkern 1.0mu\tau_{n}f}\right]$ is the frequency correlation function. Assuming a uniform planewave normal distribution, $\mathbf{r}_{n}=\begin{bmatrix}\cos{\left(\phi_{n}\right)}&\sin{\left(\phi_{n}\right)}&0\end{bmatrix}^{\mathsf{T}}$ with $\phi_{n}\sim\mathcal{U}\left(0,2\pi\right)$, as in Clarke’s two-dimensional model [10, Sec. 5.4], and a uniform excess delay distribution $\tau_{n}\sim\mathcal{U}\left(0,\tau_{\text{max}}\right)$, the covariance functions are given as
$c_{\mathbf{x}}\left(\mathbf{x}\right)$ $=J_{0}\left(2\pi x\lambda^{-1}\right)$ (7)
$c_{f}\left(f\right)$ $=\frac{1-\mathrm{e}^{-\mathrm{i}\mkern 1.0mu\tau_{\text{max}}f}}{\mathrm{i}\mkern 1.0mu\tau_{\text{max}}f}$ (8)
where $x=\norm{\mathbf{x}}_{2}$ is the distance and $\lambda\in\mathbb{R}$ is the wavelength. It can observed that samples are approximately uncorrelated, $c_{\mathbf{x}},c_{f}<0.33$, for $x\geq 0.71\lambda$ and $f\geq 4.6\,\tau_{\text{max}}^{-1}$. In setting $\tau_{\text{max}}=10dc^{-1}$ with $c$ being the speed of light in a vacuum and $d$ being the distance between transmitter and receiver, the channel is approximately uncorrelated for a frequency shift of $f=1.4$ MHz at a distance of $100$ m, and approximately uncorrelated for a frequency shift of $f=6.9$ MHz at a distance of $20$ m. Thus, given sufficient bandwidth and distance, sampling across frequency and space is equivalent of drawing independent random samples from the same distribution. The usage of frequency as a proxy for space, is further verified through the DT in Sec. III.
III Scenario
@@
Evaluation of the proposed system is done using the data from the measurement campaign described in [11]. The data consists of wideband channel measurements done with a stationary receiver and 127 different transmitter locations. Both the transmitter and receiver are equipped with omnidirectional antennas. The layout of the measurement campaign is illustrated on Fig. 1 Left. The layout of the digital twin of the same scenario is illustrated on Fig. 1 Right, with the digital-twin (DT) equivalent 127 transmitter locations marked as green dots. At each of the 127 locations, 8001 equispaced frequency domain channel measurements are made, spanning the bandwidth from 2 GHz to 10 GHz, yielding a frequency resolution of 1 MHz. The geometry used in the digital twin is extracted from OpenStreetMap [7] using the Blosm for Blender blender plugin and imported into the sionna python library [12]. All surfaces, ground included, are assigned the sionna-defined itu_brick material which has a relative permittivity of $\epsilon_{r}=3.91$ [13, table 3].
III-A Approximating the statistics of small-scale fading via frequency-domain sampling
(a) Near the transmitter
(b) Far from the transmitter

### v2

# Prediction of Wireless Channel Statistics with Ray Tracing and Uncalibrated Digital Twin

@@ Abstract
We introduce a framework for predicting wireless channel statistics based on digital twin (DT) and ray tracing. The DT is derived from satellite images and is uncalibrated, as it does not assume precise information on the electromagnetic properties of the materials in the environment. The uncalibrated DT is utilized to derive a geometric prior that informs a Gaussian process (GP) and thereby predict channel statistics using only a few measurements. The framework also quantifies uncertainty, offering statistical guarantees for rate selection in ultra-reliable low-latency communication (URLLC). Experimental validation demonstrates the efficacy of the proposed framework using measurement data.
@@ I Introduction
Digital twins (DTs) that accurately model wireless channels have demonstrated significant utility for both communication and localization tasks [1]. For instance, by maintaining a precise DT, site-specific digital twins can be used to reduce CSI acquisition and feedback overhead [2]. However, for the DT to provide accurate predictions, its electromagnetic properties should match real-world channel conditions.
@@
The process of configuring the objects and materials in the DT to represent the wireless environment is referred to as calibration. Differentiable ray tracers (RT), such as Sionna RT [3] are instrumental in enabling effective calibration methods. Differentiable RTs allow the computation of gradients for system parameters, enabling tasks like data-driven material property optimization [4].
This work addresses resource allocation for ultra-reliable low-latency communication (URLLC) communications, aiming to select a data rate that meets a statistical reliability constraint. Due to latency limitations, channel feedback may be infeasible and hence the system must rely on alternatives such as channel statistics maps to select a rate which satisfies the reliability constraint with a sufficiently high probability. While a fully calibrated DT can directly predict channel statistics by accurately modeling objects and materials, frequent calibration is resource-intensive and limits its practical use. To address this, we focus on exploiting uncalibrated DTs that can be directly derived from satellite images available through open-source maps like OpenStreetMap [5], without relying on prior knowledge of material properties in the scene or requiring data-intensive calibration processes.
@@
We summarize our contributions and results as follows. An uncalibrated DT is used to extract geometric features of the environment. These are combined with channel measurements from a few anchor points to estimate the channel statistics across the entire scene. The estimation leverages the spatial interpolation capabilities of a Gaussian process (GP) along with the geometric consistency of the DT to predict channel statistics with high accuracy in the region of interest. This is validated via experimental data obtained through a measurement campaign. The results show that the proposed geometry-informed GP significantly improves prediction accuracy across the entire scene compared to the benchmarks where (i) an uncalibrated DT directly predicts channel statistics; and (ii) a GP uses location information to predict channel statistics. Moreover, when the channel statistics map is used in rate selection for URLLC, the proposed framework is shown to, approximately, double the data rate while maintaining the same reliability, compared to the benchmarks. In summary, the geometry-informed GP framework scales the small number of calibration measurements to channel predictions that are statistically accurate for the entire region of interest.
Organization
Section II presents the problem formulation, and the construction of the uncalibrated DT. Section III presents the channel-prediction framework, and Section IV presents an effective method for experimentally measuring channel-fading statistics. Finally, Section V details the rate-selection framework and its experimental validation.
II System model and Problem Formulation
We consider a system where an access point (AP) serves a user (UE) located at position $\mathbf{x}$. The UE, equipped with a single antenna, acts as the transmitter (TX) and transmits its zero-mean, unit-power symbol $s\in\mathbb{C}$ with transmit power $P_{\text{tx}}\geq 0$. The single-antenna AP then functions as the receiver (RX). The received narrowband signal, after delay-spread equalization, is
$y=\sqrt{P_{\text{tx}}}h\left(\mathbf{x}\right)s+n,$ (1)
@@ II System model and Problem Formulation
where $h\left(\mathbf{x}\right)\in\mathbb{C}$ is the channel between the transmitter and receiver, and $n\sim\mathcal{CN}\left(0,\sigma_{n}^{2}\right)$ is additive white Gaussian noise (AWGN). Thus, the SNR given by $\gamma\left(\mathbf{x}\right)=\frac{P_{\text{tx}}\absolutevalue{h\left(\mathbf{x}\right)}^{2}}{\sigma_{n}^{2}}$.
II-A Problem formulation
For a user with position $\mathbf{x}$, the goal is to select the maximum rate $R(\mathbf{x})$ such that the outage probability is at most $\epsilon\in(0,1)$. The outage probability represents the probability of exceeding the channel capacity, $p_{\text{out},\mathbf{x}}=P\big{(}{\log_{2}({1+\gamma({\mathbf{x}}}))<R(\mathbf{x})}\big{)}$. If the distribution of $h\left(\mathbf{x}\right)$ is known, then $R(\mathbf{x})$ can be selected using the statistics of the channel power $p(\mathbf{x})=|h\left(\mathbf{x}\right)|^{2}$. However, due to latency constraints and the dynamic properties of the wireless communication channel, the distribution of $p(\mathbf{x})$ is not perfectly known. To account for the uncertainty associated with $p_{\text{out},\mathbf{x}}$, this work aims to maximize the rate $R(\mathbf{x})$ that can be selected while ensuring compliance with a reliability constraint that bounds the meta-probability as $\tilde{p}_{\epsilon}\overset{\Delta}{=}P\left(p_{\text{out,$\mathbf{x}$}}>\epsilon\right)\leq\delta$. This can also be expressed as [6]
@@ II-A Problem formulation
$\tilde{p}_{\epsilon}=P\left(R(\mathbf{x})>\log_{2}\left(1+\frac{P_{\text{tx}}p_{\epsilon}\left(\mathbf{x}\right)}{\sigma_{n}^{2}}\right)\right)\leq\delta.$ (2)
Here, $p_{\epsilon}\left(\mathbf{x}\right)$ is the $\epsilon$-quantile the channel power. In this work, we rely on statistical radio maps to estimate the power quantiles $p_{\epsilon}\left(\mathbf{x}\right)$. To this end, we employ an uncalibrated digital twin (DT) as a geometric prior to inform a Gaussian process (GP), enabling channel-statistic predictions from only a few measurements.
Traditional radio maps require dense measurements and lack geometric context. On the other hand, while capturing the underlying geometry, uncalibrated digital twins yield inaccurate channel statistics, as ray tracing relies on assumed electromagnetic properties that may not reflect real materials. This mismatch distorts the received power, impacting the prediction of URLLC-relevant statistics (i.e., lower quantiles). To address these limitations, our proposed framework, presented in Section III, leverages geometry-based DT outputs to embed spatial correlations within a GP, which utilizes a small set of measurements to accurately predict the channel statistics. GPs are employed here for their sample-efficient spatial interpolation properties, enabling not only efficient prediction but also closed-form uncertainty quantification [7].
Subsequently, in contrast to conventional radio maps that provide point estimates of average statistics, the proposed framework provides predictions of URLLC-relevant statistics along with closed-form quantification of prediction uncertainty. This enables the selection of transmission rates under statistical guarantees. In Section V, we validate our GP-based framework using real measurements, demonstrating its ability to accurately predict channel statistics and meet URLLC reliability targets.
@@
Figure 1: Bottom right: A satellite picture of the outdoor scenario not taken at the day of the measurement. Transmitter locations marked with blue dots [6]. Top left: Illustration of the geometry made using the sionna.rt.scene.render python library [3] and data extracted from OpenStreetMap [5]. Transmitter locations are marked as green dots.
II-B Scene Description and Digital‐Twin Formation
@@
We consider the scenario illustrated in Fig. 1 (Left), featuring a stationary receiver and 127 distinct transmitter locations, each equipped with an omnidirectional antenna. A digital twin (DT) of this environment was created by extracting the geometry from OpenStreetMap[5] using the Blosm for Blender plugin, and importing it into the Sionna Python library[3]. In the DT (Fig. 1 (Right)), the 127 transmitter sites are marked as green dots. All surfaces, including the ground, are modeled with Sionna’s predefined itu_brick material (relative permittivity $\epsilon_{r}=3.91$) [8, Table 3]. No calibration of the DT is performed.
III Site-specific Channel Statistics Prediction
In this section, we focus on channel statistics prediction, with the rate-selection problem addressed and validated using experimental data in Section V-C.
III-A Data Preprocessing and Estimation of Quantiles
Consider an input dataset $\mathcal{\hat{D}}=\{\mathbf{p}_{d},\mathbf{x}_{d}\}_{d=1}^{D}$, where for each location $\mathbf{x}_{d}$, $\mathbf{p}_{d}=\left(p_{d,1},\ \ldots,\ p_{d,N}\right)$ captures $N$ independent small-scale fading channel measurements. To enable spatial channel prediction, the logarithm of the $\epsilon$-quantile channel power, defined as $q_{\epsilon}\left(\mathbf{x}\right)=\ln\left(p_{\epsilon}\left(\mathbf{x}\right)\right)$, is modeled as a Gaussian process (GP). The logarithmic transformation is applied to allow transformed data to be effectively modeled as a GP. To estimate the log $\epsilon$-quantile $q_{\epsilon}(\mathbf{x})$, the log-fading power quantile at a location $\mathbf{x}_{d}$ is estimated as $\widehat{q}_{\epsilon,d}=\ln\left({p}_{d,(r)}\right)$ for $r=\lfloor N\epsilon\rfloor$ and $p_{d,(r)}$ denoting the $r$-th order statistic of $\mathbf{p}_{d}$. Indeed, this estimate is unbiased and asymptotically behaves as a Gaussian distribution [9]. Subsequently, the input to the channel prediction scheme is a dataset $\mathcal{{D}}=\{\widehat{q}_{\epsilon,d},\mathbf{x}_{d}\}_{d=1}^{D}$.
III-B GP for Channel Statistics Prediction
By modeling the quantile function $q_{\epsilon}\left(\mathbf{x}\right)$ as a Gaussian process, GP regression can be used to predict fading statistics by interpolating channel measurements resulting in a predictive distribution. This approach was proposed in [9] to learn a predictive distribution $q_{\epsilon}^{\sf{spatial}}(\mathbf{x}|\mathcal{D})\sim\mathcal{N}(\mu(\mathbf{x}|\mathcal{D}),\sigma^{2}(\mathbf{x}|\mathcal{D})),$ where $\mu(\mathbf{x}|\mathcal{D})$ is the predictive mean and $\sigma^{2}(\mathbf{x}|\mathcal{D})$ is the predictive variance. For the implementation details, we refer the reader to [9, Sec. III-B]. Notably, in this work, we go beyond spatial interpolation by incorporating geometric correlations extracted from an uncalibrated DT to enable geometry-inspired statistical predictions.
For each position $\mathbf{x}$, a DT can be used to obtain the CDF of the received power ${\mathbf{\psi}}^{\sf{DT}}(\beta,\mathbf{x})=P(p(\mathbf{x})\leq\beta)$. If the DT is perfectly calibrated then ${\mathbf{\psi}}^{\sf{DT}}(\beta,\mathbf{x})$ can be directly used to provide an estimate $q^{\sf{direct}}_{\epsilon}(\mathbf{x})$. However, in this work, we consider uncalibrated DTs generated from open-source maps (e.g. OpenStreetMap [5]) and with default material properties. Thus, due to mismatched material properties, changes in the wireless environment and other non-idealities, the direct use of a DT does not necessarily provide an accurate estimate of $q_{\epsilon}(\mathbf{x})$. However, the DT’s received power CDF ${\mathbf{\psi}}^{\sf{DT}}(\beta,\mathbf{x})$ can be leveraged to provide site-specific geometric features.
We propose to exploit the geometric information implicitly conveyed by ${\mathbf{\psi}}^{\sf{DT}}(.)$ and a small number of channel measurements to predict fading statistics across the entire site. In particular, by processing the DT’s distribution of received power $\mathbf{\psi}^{\sf{DT}}(\beta,\mathbf{x})$, we employ the following predictive distribution
$q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$ $\sim\mathcal{N}(\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D}),\sigma_{\sf{DT}}^{2}(\mathbf{y}|\mathcal{D})),$
Here $\mathbf{f}(\cdot)$ uniformly samples $100$ points from the DT’s CDF of received power $\mathbf{\psi}^{\sf{DT}}(\beta,\mathbf{x})$.
With the $\mathbf{y}$ defined above, the GP kernel $k(\mathbf{y},\mathbf{y}^{\prime})$ not only captures spatial patterns but also implicitly encodes geometric information provided by the DT. In particular, for two receiver positions $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ that result in similar DT outputs, $\mathbf{\psi}^{\sf{DT}}(\beta,\mathbf{x}_{1})\approx\mathbf{\psi}^{\sf{DT}}(\beta,\mathbf{x}_{2})$, the GP’s covariance function ensures that the predicted channel statistics at these two locations are similar. Thus, the DT’s geometric features are captured via the GP’s covariance function, embedding geometric correlations that allows efficient prediction of channel statistics using a limited number of measurements. Moreover, this proposed prediction framework is shown to provide significant improvements in prediction errors compared to the following benchmarks:
Before presenting the experimental results, which corroborate the efficacy of our framework, the next section provides theoretical and experimental justification for an effective method for the experimental characterization of small-scale fading statistics.
IV Experimental Characterization of Small-Scale Fading Statistics
Obtaining spatial samples of the power quantiles $p_{\epsilon}\left(\mathbf{x}\right)$ requires either relying on model based approaches which are susceptible to model mismatch errors, or a very high number of channel measurements to establish an empirical distribution, which results in high latency. An alternative approach is to use a range of subcarriers and use frequency as a proxy for space [6]. By measuring a wide bandwidth, an empirical channel power distribution can be established and an estimate of the $\epsilon$-quantile of the channel power, $p_{\epsilon}\left(\mathbf{x}\right)$, can be extracted.
IV-A Theoretical Analysis
The validity of using frequency as a proxy for space can be argued by considering the WSSUS Rayleigh fading channel [10]. In the farfield of the transmitter, the impinging field can be written as a position dependent distribution of planewaves. In quantizing the distribution, the received signal in the vicinity of a point $\boldsymbol{x}$ can be written as $h\left(\mathbf{x},f\right)\approx\sum_{n=1}^{N}\alpha_{n}\mathrm{e}^{-\mathrm{i}\mkern 1.0muk\mathbf{x}^{\mathsf{T}}\mathbf{r}_{n}}\mathrm{e}^{-\mathrm{i}\mkern 1.0mu\tau_{n}f}$, where $k\in\mathbb{R}$ is the wavenumber, $\mathbf{x}\in\mathbb{R}^{3\times 1}$ is a point in the vicinity of $\boldsymbol{x}$ and $\alpha_{n}\in\mathbb{C}$ is the amplitude and phase, $\mathbf{r}_{n}\in\left\{\mathbb{R}^{3\times 1}|\mathbf{r}_{n}^{\mathsf{T}}\mathbf{r}_{n}=1\right\}$ is the normal vector indicating the direction of propagation, and $\tau_{n}$ is the excess delay of the $n$’th planewave. In the Rayleigh channel, $\alpha_{n}$ is a set of zero mean independent identically distributed circular symmetric complex random variables. Independence between all $\tau_{n}$, $\mathbf{r}_{n}$, and $\alpha_{n}$ is assumed. Due to $\alpha_{n}$ being circular symmetric complex random variables, $\mathbb{E}[h\left(\mathbf{x},f\right)]=0$. Additionally, due to the central limit theorem, $h\left(\mathbf{x},f\right)\sim\mathcal{CN}\left(0,\sigma^{2}\right)$ for high $N$, where $\sigma^{2}=\mathbb{E}\left[\absolutevalue{h\left(\mathbf{x},f\right)}^{2}\right]$. To allow relating the power quantile obtained from sampling across frequency to the power quantile obtained from sampling across space, the covariance properties of the channel across frequency and space, is analyzed. The channel covariance function is given by $\Sigma^{2}\left(\mathbf{x}_{1},\mathbf{x}_{2},f_{1},f_{2}\right)=\mathbb{E}\left[h\left(\mathbf{x}_{1},f_{1}\right)h^{*}\left(\mathbf{x}_{2},f_{2}\right)\right],=\absolutevalue{s}^{2}\sigma^{2}c_{\mathbf{x}}\left(\mathbf{x}_{1}-\mathbf{x}_{2}\right)c_{f}\left(f_{1}-f_{2}\right),$ where $c_{\mathbf{x}}\left(\mathbf{x}\right)=\mathbb{E}\left[\mathrm{e}^{-\mathrm{i}\mkern 1.0muk\mathbf{x}^{\mathsf{T}}\mathbf{r}_{n}}\right]$ is the spatial correlation function, and $c_{f}\left(f\right)=\mathbb{E}\left[\mathrm{e}^{-\mathrm{i}\mkern 1.0mu\tau_{n}f}\right]$ is the frequency correlation function. Assuming a uniform planewave normal distribution, $\mathbf{r}_{n}=\begin{bmatrix}\cos{\left(\phi_{n}\right)}&\sin{\left(\phi_{n}\right)}&0\end{bmatrix}^{\mathsf{T}}$ with $\phi_{n}\sim\mathcal{U}\left(0,2\pi\right)$, as in Clarke’s two-dimensional model [10, Sec. 5.4], and a uniform excess delay distribution $\tau_{n}\sim\mathcal{U}\left(0,\tau_{\text{max}}\right)$, the covariance functions are given as $c_{\mathbf{x}}\left(\mathbf{x}\right)=J_{0}\left(2\pi x\lambda^{-1}\right)$ and $c_{f}\left(f\right)=\frac{1-\mathrm{e}^{-\mathrm{i}\mkern 1.0mu\tau_{\text{max}}f}}{\mathrm{i}\mkern 1.0mu\tau_{\text{max}}f}$, where $x=\norm{\mathbf{x}}_{2}$ is the distance and $\lambda\in\mathbb{R}$ is the wavelength. It can observed that samples are approximately uncorrelated, $c_{\mathbf{x}},c_{f}<0.33$, for $x\geq 0.71\lambda$ and $f\geq 4.6\,\tau_{\text{max}}^{-1}$. In setting $\tau_{\text{max}}=10dc^{-1}$ with $c$ being the speed of light in a vacuum and $d$ being the distance between transmitter and receiver, the channel is approximately uncorrelated for a frequency shift of $f=1.4$ MHz at a distance of $100$ m, and approximately uncorrelated for a frequency shift of $f=6.9$ MHz at a distance of $20$ m. Thus, given sufficient bandwidth and distance, sampling across frequency and space is equivalent of drawing independent random samples from the same distribution.
IV-B Numerical Validation